# Ariadne: A Controllable Framework for Probing and Extending VLM Reasoning Boundaries

### v1

@@ Abstract
While Vision-Language Models (VLMs) post-trained with Reinforcement Learning (RL) show impressive general reasoning, their evaluation is often confined to language-dominant tasks (e.g., math). This raises a critical question: can RL post-training truly extend the inherent capability boundary of a base VLM, particularly for visual-centric spatial tasks where it initially fails? To investigate this, we introduce Ariadne, a framework utilizing synthetic mazes for multi-step spatial reasoning where task difficulty (e.g., path length, turns) is precisely controlled. We leverage this controllable environment to train VLMs using Reinforcement Learning with Verified Rewards (RLVR) in a difficulty-aware curriculum. Surprisingly, post-RLVR training, the VLM achieves over 50% accuracy on a problem set where the base model scored 0%, demonstrating that our approach expands the model’s initial capability boundary. To assess real-world viability, we evaluate out-of-distribution (OOD) generalization on practical benchmarks. Despite training only on synthetic maze samples, Ariadne achieves significant zero-shot improvements, averaging 16% on MapBench (e.g., museum navigation) and 24% on ReasonMap (subway transfer tasks). These results confirm that our method not only broadens the model’s fundamental limits but also enhances its generalization to real-world spatial reasoning. We acknowledge our study is limited to the post-training phase, given the opaqueness of pre-training data, and hope our research motivates further work on specialized, capability-extending alignment.
@@ 1 Introduction
  • We propose Ariadne, a controllable framework to explore VLM reasoning boundaries post-RLVR training using purely synthetic maze tasks.
  • Ariadne-trained VLMs achieve over 50% accuracy on maze problems where the base model scored 0%, demonstrating a clear expansion of the initial capability boundary.
  • Beyond synthetic tasks, Ariadne achieves notable zero-shot improvements on real-world benchmarks (MapBench [10] and ReasonMap [11]), confirming that RLVR can broaden VLM reasoning across both in-domain and out-of-distribution settings.
@@ 2.1 Reasoning in VLMs
VLMs have achieved substantial progress on tasks such as Visual Question Answering (VQA) and image captioning [12, 13, 14]. To strengthen reasoning capabilities, prior work has explored Chain of Thought (CoT) prompting [15] and the construction of Supervised Fine-Tuning (SFT) datasets enriched with step-level rationales [14]. While these strategies enhance reasoning to some extent, they fall short of capturing human-like cognitive processes such as questioning and self-verification, and their effectiveness diminishes on complex reasoning tasks. Recent models such as DeepSeek-R1 [4] attempt to address this gap by combining cold-start initialization with RL, thereby acquiring higher-quality multimodal CoT reasoning and achieving state-of-the-art results on challenging visual reasoning benchmarks.
@@ 4.3.1 RLVR Enables VLMs to Break Reasoning Boundaries
Figure 3 presents the reward trajectory during GRPO training on AlphaMaze, along with quantitative evaluations on the test set. Across all evaluated movement step ranges, Ariadne shows a consistent advantage over its base model, Qwen2.5-VL-7B-Instruct. For Qwen2.5-VL-7B-Instruct, we perform eight rollouts to confirm model stability and observe that performance drops to zero when the path requires three movement steps or three turns, which we define as the model’s reasoning boundary. After RLVR training, Ariadne’s success rate rises to over 50% on 3 step cases and over 10% on 3 turn cases, and the collapse point shifts from 3 to 5, indicating that RLVR effectively extends the model’s reasoning boundary to more complex path configurations.
The qualitative examples in Figure 4 further illustrate these results. Successful trajectories generally occur in mazes with simple layouts and moderate detour requirements, allowing coherent path following. In contrast, failure cases are often associated with local structural complexity, dense turns, narrow corridors, and elongated detours that make it difficult for the model to maintain consistent global navigation.
@@ 4.3.2 RLVR Extends VLM Reasoning to Real-world Tasks
As shown in Figure 5, the two models exhibit a clear divergence in real-world map-based navigation across diverse task settings. The baseline Qwen2.5-VL-7B-Instruct exhibits localization and planning errors, such as misidentifying targets, taking inefficient detours, or violating environmental constraints like rivers and walls, leading to incomplete or infeasible trajectories. In contrast, Ariadne employs fine-grained, node-level planning that maintains goal-directed navigation while respecting structural boundaries. This behavior generalizes across both unstructured outdoor networks (e.g., trails) and structured indoor layouts (e.g., museums), demonstrating the effectiveness of GRPO adaptations beyond synthetic mazes.
@@ 5 Conclusion
In this work, we introduced Ariadne, a controllable framework for systematically probing and extending the visual reasoning capabilities of VLMs through RLVR. Our experiments demonstrate that RLVR effectively extends the reasoning boundaries of VLMs, enabling models trained on short and simple trajectories to generalize to longer and more complex cases. Furthermore, the improvements transfer robustly to real-world settings, even under visual domain shifts, indicating that visual reasoning skills acquired in synthetic environments are transferable across diverse visual contexts. However, we observe a sharp decline in performance beyond a critical complexity threshold. This limitation suggests that RLVR may only strengthen and extend reasoning patterns learned during pretraining, rather than inducing entirely new reasoning mechanisms absent from the pretraining data. Future work will explore how controllable reinforcement learning can be integrated into the pretraining stage to further expand the fundamental reasoning capacity of VLMs.

### v2

@@ Abstract
While Vision-Language Models (VLMs) post-trained with Reinforcement Learning (RL) show impressive general reasoning, their evaluation is often confined to language-dominant tasks (e.g., math). This raises a critical question: can RL post-training truly extend the inherent capability boundary of a base VLM, particularly for visual-centric spatial tasks where it initially fails? To investigate this, we introduce Ariadne, a framework utilizing synthetic mazes for multi-step spatial reasoning with precisely controlled task difficulty (e.g., path length, turns). We employ this controllable setup to train VLMs using Reinforcement Learning with Verified Rewards (RLVR) under a difficulty-aware curriculum. Surprisingly, after RLVR training, the VLM attains over 50% accuracy on tasks where the base model scored 0%, indicating that RLVR can extend the base policy’s capability boundary within, but not fully beyond, the training data scope. Specifically, it generalizes to unseen numbers of turns but fails when the number of movement steps exceeds training exposure, revealing a divergent phenomenon in spatial reasoning. Furthermore, we observe that in real-world settings, this boundary shifts: the model handles longer-step reasoning, likely due to noise-induced redundancy in natural scenarios. Despite being trained solely on synthetic mazes, Ariadne still yields notable zero-shot gains, averaging 16% on MapBench (e.g., museum navigation) and 24% on ReasonMap (subway transfer tasks). These results suggest that our method not only broadens the model’s intrinsic limits but also enhances its generalization to realistic spatial reasoning tasks. We note that our study is confined to the post-training phase due to the opacity of pre-training data and hope this work inspires further exploration into specialized, capability-extending alignment.
@@ 1 Introduction
  • We propose Ariadne, a controllable framework for probing and extending VLM reasoning boundaries via RLVR on synthetic maze tasks with precisely tunable difficulty.
  • Ariadne achieves over 50% accuracy on maze problems where the base model scored 0%, indicating that RLVR can extend the base policy’s capability boundary within but not fully beyond the training domain, generalizing to unseen numbers of turns but failing on longer step sequences.
  • We identify a divergent phenomenon where real-world reasoning boundaries deviate from synthetic ones, as models generalize better to longer-step reasoning in noisy, natural environments.
  • Despite being trained solely on synthetic mazes, Ariadne delivers consistent zero-shot gains on real-world benchmarks, MapBench [10] (16%) and ReasonMap [11] (24%), highlighting its effectiveness in enhancing spatial reasoning beyond training conditions.
@@ 2.1 Reasoning in VLMs
VLMs have achieved substantial progress on tasks such as Visual Question Answering (VQA) and image captioning [12, 13, 14]. To strengthen reasoning capabilities, prior work has explored Chain of Thought (CoT) prompting [15] and the construction of Supervised Fine-Tuning (SFT) datasets enriched with step-level rationales [14]. While these strategies enhance reasoning to some extent, they fall short of capturing human-like cognitive processes such as questioning and self-verification, and their effectiveness diminishes on complex reasoning tasks. Recent models such as DeepSeek-R1 [4] attempts to address this gap by combining cold-start initialization with RL, thereby acquiring higher-quality multimodal CoT reasoning and achieving state-of-the-art results on challenging visual reasoning benchmarks.
@@ 4.3.1 RLVR Enables VLMs to Break Reasoning Boundaries
Figure 3 presents the reward trajectory during GRPO training on AlphaMaze, along with quantitative evaluations on the test set. Across all evaluated movement step ranges, Ariadne shows a consistent advantage over its base model, Qwen2.5-VL-7B-Instruct [18]. For Qwen2.5-VL-7B-Instruct, we perform eight rollouts to confirm model stability and observe that performance drops to zero when the path requires three movement steps or three turns, which we define as the model’s initial reasoning boundary. After RLVR training, Ariadne’s success rate rises to over 50% on 3-step cases and over 10% on 3-turn cases, with the collapse point shifting from 3 to 5. This shows that RLVR extends the reasoning boundary partially: the model generalizes to unseen numbers of turns but struggles when the number of movement steps exceeds its training range. We term this a partially extended reasoning boundary and a divergent generalization phenomenon, indicating that the improvement occurs along specific dimensions rather than uniformly across spatial factors.
@@
The qualitative examples in Figure 4 further illustrate these dynamics. Successful trajectories typically occur in mazes with simple layouts and moderate detours, where spatial coherence can be preserved. In contrast, failure cases are concentrated in locally complex regions with dense turns, narrow corridors, or elongated detours, which disrupt global navigation consistency and highlight the limits of Ariadne’s extended but still bounded spatial reasoning capability.
@@ 4.3.2 RLVR Extends VLM Reasoning to Real-world Tasks
Figure 7: Distribution of successful trajectory step counts generated by Ariadne and Qwen2.5-VL-7B-Instruct on the MapBench benchmark. Ariadne is trained exclusively on trajectories with 1 to 5 steps. This range is marked as ’In-Distribution (Step Counts)’ (light green). The ’Out-of-Distribution (Step Counts)’ region (dark green) evaluates the models’ ability to generalize to longer, unseen path lengths (6 steps and beyond).
As shown in Figure 5, the two models exhibit a clear divergence in real-world map-based navigation across diverse task settings. The baseline Qwen2.5-VL-7B-Instruct [18] exhibits localization and planning errors, such as misidentifying targets, taking inefficient detours, or violating environmental constraints like rivers and walls, leading to incomplete or infeasible trajectories. In contrast, Ariadne employs fine-grained, node-level planning that maintains goal-directed navigation while respecting structural boundaries. This behavior generalizes across both unstructured outdoor networks (e.g., trails) and structured indoor layouts (e.g., museums), demonstrating the effectiveness of GRPO adaptations beyond synthetic mazes.
@@
Figure 7 further examines this phenomenon on the MapBench benchmark. Ariadne, trained only on trajectories with 1–5 steps, maintains strong performance within this in-distribution range while showing selective generalization to out-of-distribution paths beyond six steps. Although the model’s training boundary is constrained to short-step trajectories, it still achieves nontrivial success on longer, unseen paths, indicating a partially extended reasoning boundary. This divergent generalization suggests that real-world spatial reasoning may be influenced by noisy or redundant structures, causing the effective reasoning boundary to become misaligned with that observed in controlled synthetic settings.
@@ 5 Conclusion
In this work, we introduced Ariadne, a controllable framework for systematically probing and extending the visual reasoning capabilities of VLMs through RLVR. Our experiments show that RLVR can effectively extend the reasoning boundary of VLMs, allowing models trained on short and simple trajectories to generalize to longer and more complex cases. However, this extension is partial and dimension-specific: the model generalizes well to unseen numbers of turns but fails when the number of movement steps exceeds its training range. This behavior represents a divergent generalization phenomenon, where reasoning improvements emerge unevenly across task dimensions rather than through a uniform expansion of capability. We further find that the reasoning boundary observed in real-world scenarios does not align with that in synthetic environments; models perform well on OOD step-count reasoning under noisy or redundant conditions, suggesting that environmental complexity may shift the effective boundary of spatial reasoning. Despite these nuances, the learned improvements transfer robustly to real-world settings, even under visual domain shifts, demonstrating that visual reasoning skills acquired in synthetic environments remain transferable across diverse contexts. Nonetheless, performance declines sharply beyond a critical complexity threshold, implying that RLVR primarily strengthens and reorganizes reasoning patterns inherited from pretraining rather than creating entirely new mechanisms. Future work will investigate how controllable reinforcement learning can be integrated into the pretraining stage to further expand the fundamental reasoning capacity of VLMs.